OpenStudy (anonymous):

Find the distance between these two points be sure that your answer is simplified and in radical form. (0,-6) and (3,4)

OpenStudy (anonymous):

from 0 to 3 is 3 units from -6 to 4 is 10 units. pythagoras gives \[d^2=3^2+10^2=109\] so \[d=\sqrt{109}\]

OpenStudy (anonymous):

thanks you, thats what i got to

OpenStudy (anonymous):

yw